# Long-Lived Photon Blockade with Weak Optical Nonlinearity

### v2

@@ Abstract
In conventional photon blockade, the occupation of a cavity mode by more than one photon is suppressed via strong optical nonlinearity. An alternative, called unconventional photon blockade, can occur under weak nonlinearity by relying on quantum interference between fine-tuned cavities. A serious limitation is the very short antibunching time window, orders of magnitude less than the cavity lifetime. We present a method to achieve photon blockade over a large time window of several cavity lifetimes, even exceeding that of conventional photon blockade, while still requiring only weak nonlinearity. This “long-lived photon blockade” (LLPB) occurs when the single-photon Green’s function exhibits a zero at a large cavity loss rate, which is satisfied by an exemplary configuration of four coupled cavities under weak driving. Our analytical results agree well with wavefunction Monte Carlo simulations. The LLPB phenomenon may aid the development of single-photon sources utilizing materials with weak optical nonlinearities.
Introduction—Photon blockade is a quantum effect whereby one photon in a nonlinear resonator blocks the entry of other photons, giving rise to antibunched photon statistics [1, 2, 3, 4]. Conventionally, photon blockade requires the optical nonlinearity within a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate. This can be interpreted in terms of a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Achieving this in the optical regime (including telecom wavelengths) is challenging, as material nonlinearities tend to be very weak relative to cavity linewidths. Systems featuring exceptionally strong optical nonlinearities are required, such as exciton-polaritons [25, 26], trions [27], Rydberg excitons [28], and Moire excitons [29].
@@
$J\gg\gamma,$ (1)
which seems to make intuitive sense: for interference to play a role, the photons must be able to hop many times before leaking away [30]. In this regime, it then turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can produce photon antibunching. However, even if the second-order photon correlation $g^{(2)}(\tau)$ vanishes at $\tau=0$, the suppression only holds over a time window of
@@
due to the Rabi-like oscillations that $J$ induces in the time-dependent state amplitudes. This is much smaller than the antibunching time window of $\delta\tau\sim 1/\gamma$ for conventional photon blockade, and poses a serious practical obstacle to the realization and exploitation of UPB. For example, a proposal for UPB at telecom wavelengths [38] requires $J\approx 20\gamma$ and yields an antibunching time window of $\sim 100\,\text{ps}$, around the resolution limit for current photodetectors. In such two-cavity models, the UPB condition reduces to $\gamma^{3}\propto\alpha J^{2}$, implying that the cavity quality factors satisfy $Q^{3}\propto\omega_{0}^{3}\delta\tau^{2}/\alpha$, where $\omega_{0}$ is the cavity resonance frequency [32]. Given a lower bound on $\delta\tau$ imposed by detector resolution limits, this condition sets a minimum requirement on $Q$ that is unfavorable in the optical regime, where $\omega_{0}$ is large and optical nonlinearities are typically weak. Flayac and Savona have suggested that with one-way dissipative inter-cavity couplings, the rapid oscillations in $g^{(2)}(\tau)$ can be avoided [53]; however, such couplings are highly challenging to implement experimentally.
@@
In this paper, we show that interference-assisted photon blockade does not require the strong-coupling condition (1), even with weakly nonlinear cavities and ordinary couplings. Using a Green’s function formalism, we analyze the steady-state behavior of coupled nonlinear cavities in the weak-driving regime [32], and show that antibunching solutions appear in the vicinity of 1-photon dark states, where the 1-photon Green’s function vanishes in a specific cavity. In the two-cavity systems where UPB was first discovered [30, 31, 32], as well as more complicated coupled-cavity configurations [52], such dark states occur only at $\gamma=0$, giving rise to UPB solutions with nonzero but small $\gamma$ satisfying Eq. (1). However, certain coupled-cavity configurations host other families of dark states at large $\gamma$ (relative to $J$). We demonstrate this with a four-cavity system, showing that it produces a qualitatively different form of antibunching that we call long-lived photon blockade (LLPB). Unlike the earlier proposal involving one-way dissipative couplings [53], our four-cavity setup involves only standard (i.e., two-way and Hermitian) inter-cavity couplings. The antibunching condition $g^{(2)}(0)=0$ is satisfied alongside weak coupling ($J\ll\gamma$) and weak nonlinearity ($\alpha\ll\gamma$). Moreover, $g^{(2)}(\tau)$ is suppressed over a time window $\delta\tau\approx 8/\gamma$, two orders of magnitude wider than a comparable two-cavity system exhibiting UPB, and even larger than for conventional photon blockade under strong nonlinearity. Thus, it may be possible to achieve photon blockade in the visible regime with weakly nonlinear optical media.
@@
Model—We consider coupled nonlinear cavities, indexed by $i=1,\dots$, under a weak coherent drive. Assuming for simplicity that all the cavities have identical parameters, the cavity Hamiltonian in the frame co-rotating with the driving field, excluding losses, is
@@
Here, $\hbar=1$, $\Delta$ is the cavity detuning relative to the driving frequency, $\alpha$ is the nonlinearity strength (Kerr coefficient), $J_{ij}$ is the coupling between cavities $i$ and $j$, and $a_{i}$ is the photon annihilation operator for cavity $i$. We assume no magneto-optic activity, so $J_{ij}=J_{ji}\in\mathbb{R}$ with a suitable gauge.
We assume a single-input, single-output setup where a cavity $d$ is driven, and we look at the photon populations in a “readout” cavity (which may or may not be $d$). The driving Hamiltonian is $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$, where $F_{d}$ is the complex driving amplitude. (The following derivation can also be straightforwardly generalized to the case of multiple coherently driven cavities.) The density matrix $\rho$ is governed by the Lindblad master equation [54]
@@
$i\frac{d|\psi\rangle}{dt}$ $=\big{[}\mathcal{H}(z)+\mathcal{H}_{d}\big{]}|\psi\rangle,$ (5)
$\mathcal{H}(z)$ $\equiv\sum_{i\neq j}J_{ij}a^{\dagger}_{i}a_{j}+\sum_{i}za^{\dagger}_{i}a_{i}+\alpha\sum_{i}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i},$ (6)
@@
where the non-Hermitian effective Hamiltonian $\mathcal{H}(z)$ is derived from (3) by adding loss to each site.
We seek a solution that is stationary (i.e., co-rotating with the drive, $d/dt\rightarrow 0$). Truncating to superpositions of only 1- and 2-photon states yields the following equal-time second-order correlation function:
$g_{ij}^{(2)}$ $=\frac{\langle a^{\dagger}_{j}a^{\dagger}_{i}a_{i}a_{j}\rangle}{\langle a^{\dagger}_{j}a_{j}\rangle\langle a^{\dagger}_{i}a_{i}\rangle}=\big{|}f_{ij}(z)\big{|}^{2},$ (8)
$f_{ij}(z)$ $=\frac{\sqrt{2}\langle i,j|\bar{\psi}^{(2)}\rangle}{\langle i|\bar{\psi}^{(1)}\rangle\langle j|\bar{\psi}^{(1)}\rangle},$ (9)
where $|i\rangle=a_{i}^{\dagger}|\text{vac}\rangle$ denotes the state with one photon in cavity $i$ and $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$. The full derivation, which generalizes a previous result based on a specific lattice model [52], is given in the Supplemental Materials [55]. In Eq. (8), the expectation values are taken over the stationary state, while in Eq. (9) $|\bar{\psi}^{(1)}\rangle$ and $|\bar{\psi}^{(2)}\rangle$ denote projections of the steady-state wavefunction to the 1- and 2-photon subspaces respectively.
Using Eq. (9), we construct the following account of how photon antibunching arises. By fine-tuning the system parameters, it may be possible to satisfy $\langle i|\bar{\psi}^{(1)}\rangle=0$, which we call a 1-photon dark state, for some cavity $i$. In the absence of nonlinearity, such a condition also implies a 2-photon dark state $\langle i,j|\bar{\psi}^{(2)}\rangle=0$, since $|\bar{\psi}^{(2)}\rangle=|\bar{\psi}^{(1)}\rangle\otimes|\bar{\psi}^{(1)}\rangle/\sqrt{2}$. We find that even though the numerator and denominator of (9) both vanish, $f_{ij}(z)$ does not vanish; in fact, $g_{ij}^{(2)}=1$ for all $z$. A weak nonlinearity, however, can induce a small misalignment between the 1-photon and 2-photon dark states. Given a 1-photon dark state at $z_{0}$, the 2-photon dark state condition is shifted to a nearby value of $z_{0}+\delta z$. At this point, the numerator of (9) vanishes while the denominator is nonzero. This is our desired photon antibunching point.
@@
It is also helpful to introduce the 1-photon and non-interacting 2-photon Green’s functions,
$G(z)$ $\equiv\sum_{n}\frac{|\varphi_{n}\rangle\langle\varphi_{n}|}{z+\epsilon_{n}},$ (10)
$G^{(2)}(z)$ $\equiv\sum_{mn}\frac{|\varphi_{m},\varphi_{n}\rangle\langle\varphi_{m},\varphi_{n}|}{2z+\epsilon_{m}+\epsilon_{n}},$ (11)
where $\{|\varphi_{n}\rangle\}$ is the eigenbasis of the coupling matrix, $\left[J_{ij}\right]|\varphi_{n}\rangle=\epsilon_{n}|\varphi_{n}\rangle$, and $|\varphi_{m},\varphi_{n}\rangle\equiv|\varphi_{m}\rangle\otimes|\varphi_{n}\rangle$. Using these and assuming $\alpha$ is perturbative, Eq. (9) can be expressed as
$f_{ij}(z)$ $=1-2\alpha\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}|k,k\rangle.$ (12)
The time-delayed correlation function $g_{ij}^{(2)}(\tau)$ can be calculated by upgrading the $a_{i}$’s and $a_{i}^{\dagger}$’s to Heisenberg picture operators [55]. For $\tau>0$, we find that the non-interacting 2-photon Green’s function in Eq. (12) is simply replaced by
$G^{(2)}(z,\tau)$ $\equiv e^{-iz\tau}\sum_{mn}\frac{|\varphi_{m},\varphi_{n}\rangle\langle\varphi_{m},\varphi_{n}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (13)
This analytical form of $g_{ij}^{(2)}(\tau)$ has not previously been reported, to our knowledge. Below, we will use it to revisit UPB, and then demonstrate the LLPB phenomenon.
Conventional blockade—First, let us consider photon blockade in one cavity. In this case, we can bypass the above perturbative procedure and directly find the Green’s functions, obtaining [55]
$g_{11}^{(2)}(\tau)=\frac{1}{1+4\alpha^{2}/\gamma^{2}}+\frac{4\alpha^{2}/\gamma^{2}}{1+4\alpha^{2}/\gamma^{2}}\left(1-e^{-\frac{\gamma}{2}\tau}\right)^{2}.$ (14)
The condition $g_{11}^{(2)}(0)\approx 0$ requires $\alpha\gg\gamma$. The corresponding time window, estimated as the interval over which $g^{(2)}(\tau)<0.5$, is $\delta\tau=-4\rm{ln}(1-\sqrt{0.5})/\gamma\approx 5/\gamma$.
UPB with two cavities—For a two-cavity system, with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, the coupling matrix has eigenvalues $\epsilon_{1,2}=\pm J$ and the eigenbasis $|\varphi_{1,2}\rangle=2^{-1/2}(|1\rangle\pm|2\rangle)$. Eqs. (10)–(12) predict antibunching in the driven cavity, $g_{11}^{(2)}(0)=0$, if and only if $z^{3}\approx-\alpha J^{2}/2$ [55]. The three roots, corresponding to three 2-photon dark states, are depicted by blue circles in Fig. 1(a). They lie close to the 1-photon dark state at $z=0$, for which $\langle 1|\bar{\psi}^{(1)}\rangle\propto G_{11}=z(z^{2}-J^{2})^{-1}=0$, marked by a red cross in the figure. We focus on the root with positive cavity loss, for which
@@
Unlike the one-cavity case, this is achievable for $\alpha\ll\gamma$, but only if we enter the strong-coupling regime (1), as previously noted [30, 32]. Moreover, from Eq. (13),
$g_{11}^{(2)}(\tau)$ $\approx\left(1-e^{-\frac{\gamma}{2}\tau}\cos(\Delta\tau)\cos(J\tau)\right)^{2}.$ (16)
The fast oscillation in $\cos(J\tau)$ yields a small time window of $\delta\tau\sim 1/J$, tied to the large coupling eigenvalues ±$\pm J$. As laid out by Eqs. (8)–(12), photon antibunching occurs when the 2-photon dark state is perturbed away from the 1-photon dark state; since the latter occurs at $z=0$, the solution involves small $\gamma$ (relative to $J$). The lattice model studied in Ref. 52 also works similarly.
Long-Lived Photon Blockade—We now present a system with a 1-photon dark state at large $\gamma$. We want the linear single-photon Green’s function to satisfy $G_{id}(z)=0$ for some $i,d$, with $\gamma=-2\mathrm{Im}(z)\gg\mathrm{max}\,|J_{ij}|$, but this turns out to be impossible for $N=2$ and $N=3$ (if we relax the assumption that the cavities are identical, there can be a zero for $N=3$ if one cavity is completely lossless); for details, see the Supplemental Materials [55]. For $N=4$, however, the condition can be satisfied with the setup in Fig. 1(b), with four cavities arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$; the driven cavity is $d=1$, and photons are extracted from cavity $i=2$. This configuration can be implemented using photonic structures like the one shown in Fig. 1(c), with two ring resonators each hosting a pair of weakly-coupled clockwise and counterclockwise modes [56, 57, 58, 59]. In the Supplemental Materials, we present further details about this exemplary coupled-ring setup, including full-wave simulations showing how it can achieve the desired model parameters [55].
A Dyson series expansion in the weak-coupling regime [60] indicates that $G_{12}(z)=G_{21}(z)$ can have a zero that meets the above conditions if $J_{12}\ll J_{23},J_{34},J_{41}$. Specifically, we consider
$[J_{ij}]=\begin{pmatrix}0&J^{\prime}/k&0&J\\ J^{\prime}/k&0&J&0\\ 0&J&0&J^{\prime}\\ J&0&J^{\prime}&0\end{pmatrix},$ (17)
where $k\gg 1$. Using Eqs. (6) and (10),
$G_{21}(z)$ $=-\frac{J^{\prime}}{k}(z+z_{0})(z-z_{0})/\det\mathcal{H},$ (18)
$z_{0}$ $\equiv-iJ\sqrt{k-(J^{\prime}/J)^{2}}.$ (19)
At $z=z_{0}$, $G_{21}$ has a zero and hence $f_{22}$ has a pole (we ignore $z=-z_{0}$, which involves putting gain in the cavities). From Eqs. (7) and (19), we have $\Delta=0$ and $\gamma\approx 2\sqrt{k}J$, consistent with the weak-coupling regime.
Next, we search near $z_{0}$ for a zero of $f_{22}(z)$. Substituting Eq. (18) back into Eq. (12), we find that for $z\approx z_{0}$,
$f(z)$ $\approx\frac{(z-z_{0}-\delta z)(z-z_{0}+\delta z)}{(z-z_{0})^{2}},$ (20)
$\delta z^{2}$ $=-\dfrac{k^{2}\alpha}{2z_{0}^{2}J^{\prime 2}}\left[{\sum}_{i}G^{(2)}_{22ii}\,\big{(}G_{i1}\,\cdot\,\det\mathcal{H}\big{)}^{2}\right]_{z=z_{0}}$ (21)
where $G^{(2)}_{22ii}\equiv\langle 2,2|G^{(2)}(z)|i,i\rangle$. By further assuming $J\gg J^{\prime}$, it is approximated as [55]
$\delta z\approx\pm{\frac{\sqrt{38}}{16}}\sqrt{\alpha\gamma}e^{-i\pi/4}.$ (22)
Therefore, LLPB occurs at $z\approx z_{0}+\delta z$, as shown schematically in Fig. 1(d).
@@
In the upper panel of Fig. 1(e), we plot $g_{22}^{(2)}(0)$ versus the cavity parameters $\Delta$ and $\gamma$ using Eqs. (20)–(22). The parameters, stated in the figure caption, satisfy $\alpha\ll\gamma$ (weak nonlinearity) and $J^{\prime}/k\ll J^{\prime}\ll J\ll\gamma$ (weak coupling). The lower panel of Fig. 1(e) shows results from wavefunction Monte Carlo (WFMC) [61, 62, 63, 64] simulations that directly solve the Lindblad master equation (4), using $F_{d}=10^{-5}$ with all other parameters kept the same. Note that the WFMC simulations give $g_{22}^{(2)}(0)\ll 1$, but not exactly zero, in the weak-drive regime; the value of $g_{22}^{(2)}(0)$ increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [55].

### v3

@@ Abstract
In conventional photon blockade, the occupation of a cavity mode by more than one photon is suppressed via strong optical nonlinearity. An alternative, called unconventional photon blockade, can occur under weak nonlinearity by relying on quantum interference between fine-tuned cavities. A serious limitation is the very short antibunching time window, orders of magnitude less than the cavity lifetime. We present a method to achieve photon blockade over a large time window of several cavity lifetimes, even exceeding that of conventional photon blockade, while still requiring only weak nonlinearity. This “long-lived photon blockade” (LLPB) occurs when the single-particle Green’s function exhibits a zero at a large cavity loss rate, which is satisfied by an exemplary configuration of four coupled cavities under weak driving. Our analytical results agree well with wavefunction Monte Carlo simulations. The LLPB phenomenon may aid the development of single-photon sources utilizing materials with weak optical nonlinearities.
Introduction—Photon blockade is a quantum effect whereby one photon in a nonlinear resonator blocks the entry of other photons, giving rise to antibunched photon statistics [1, 2, 3, 4]. Conventionally, photon blockade requires the optical nonlinearity in a cavity to exceed its loss: i.e., $\alpha\gg\gamma$, where $\alpha$ characterizes the strength of the nonlinearity and $\gamma$ is the cavity decay rate. This can be interpreted as a two-photon state being driven off-resonance by the nonlinearity, shifting its energy by more than the cavity linewidth [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Achieving this in the optical regime (including telecom wavelengths) is challenging as optical nonlinearities tend to be weak relative to cavity linewidths. Systems with exceptionally strong nonlinearities are required, such as exciton-polaritons [25, 26], trions [27], Rydberg excitons [28], and Moire excitons [29].
@@
$J\gg\gamma.$ (1)
The intuition is that for interference to play a role, the photons should be able to hop many times before leaking away [30]. In this regime, it turns out that even a weak nonlinearity, $\alpha\ll\gamma$, can cause photon antibunching, i.e., the vanishing of the second-order photon correlation $g^{(2)}(\tau)$ at $\tau=0$. However, the suppression only holds over a time window
@@
due to the Rabi-like state amplitude oscillations induced by $J$. This is much smaller than the antibunching time window of $\sim 1/\gamma$ for conventional photon blockade, and poses a serious practical obstacle to the realization of UPB. For example, a proposal for UPB at telecom wavelengths [38] requires $J\approx 20\gamma$ and yields a time window of $\delta\tau\sim 100\,\text{ps}$, which is near the resolution limit of current photodetectors. In two-cavity systems, the UPB condition can be reduced to $\gamma^{3}\propto\alpha J^{2}$, and the required cavity quality factors are $Q^{3}\propto\omega_{0}^{3}\delta\tau^{2}/\alpha$, where $\omega_{0}$ is the cavity resonance frequency [32]. With a lower bound on $\delta\tau$ imposed by detector resolution limits, this sets a minimum requirement on $Q$ that is unfavorable in the optical regime, where $\omega_{0}$ is large and optical nonlinearities are typically weak. Flayac and Savona have suggested that one-way dissipative inter-cavity couplings can suppress the rapid oscillations in $g^{(2)}(\tau)$ [53], but such couplings are very challenging to implement experimentally.
@@
In this paper, we show that interference-assisted photon blockade does not inherently require the strong-coupling condition (1), even if we use ordinary couplings. In systems of weakly-nonlinear and weakly-driven cavities [32], photon blockade emerges perturbatively from linear limit solutions of the system that we call “single-particle dark state” (SPDS), defined by the vanishing of the single-particle Green’s function in a designated signal cavity. In the two-cavity systems where UPB was first discovered [30, 31, 32], as well as other cavity configurations [52], it so happens that the SPDS only exists under zero loss ($\gamma=0$), so the antibunching solutions generated perturbatively from them satisfy Eq. (1). However, the intuition behind Eq. (1) is not reliable, as there is no general principle that forbids destructive interference from occurring in lossy systems. As an example, we design a system of four lossy cavities ($\gamma\gg J$) hosting an SPDS, via the exact cancellation of multiple strongly-attenuated interference pathways. In the weakly-nonlinear quantum regime ($\alpha\ll\gamma$), this gives rise to a qualitatively different form of antibunching we call long-lived photon blockade (LLPB). The two-photon correlation function $g^{(2)}(\tau)$ is suppressed over a time window $\delta\tau\approx 8/\gamma$, two orders of magnitude wider than a comparable two-cavity system exhibiting UPB, and even larger than for conventional photon blockade under strong nonlinearity. Notably, our setup uses only couplings of the standard (reciprocal and Hermitian) type. Using these ideas, it may be possible to achieve photon blockade using photonic devices with weak optical nonlinearities.
@@
Model—Consider a general system of coupled nonlinear cavities, indexed by $i=1,\dots$, under a weak coherent drive. Assuming for simplicity that all the cavities have identical parameters, the cavity Hamiltonian in the frame co-rotating with the driving field, excluding losses, is
@@
Here, $\hbar=1$, $\Delta$ is the cavity detuning relative to the driving frequency, $\alpha$ is the nonlinearity strength (Kerr coefficient), $J_{ij}$ is the coupling between cavities $i$ and $j$, and $a_{i}$ is the photon annihilation operator for cavity $i$. There is no magneto-optic activity, so $J_{ij}=J_{ji}\in\mathbb{R}$ in a suitable gauge.
We assume a single-input, single-output setup; a single cavity $d$ is driven, with driving Hamiltonian $\mathcal{H}_{d}=F_{d}a_{d}^{\dagger}+F^{*}_{d}a_{d}$ where $F_{d}$ is the complex driving amplitude. (The following derivation can be straightforwardly generalized to multiple coherently driven cavities.) The density matrix $\rho$ is governed by the Lindblad equation [54]
@@
$i\frac{d|\psi\rangle}{dt}$ $=\big{[}\mathcal{H}+\mathcal{H}_{d}\big{]}|\psi\rangle,$ (5)
$\mathcal{H}$ $\equiv\sum_{i\neq j}J_{ij}a^{\dagger}_{i}a_{j}+\sum_{i}za^{\dagger}_{i}a_{i}+\alpha\sum_{i}a^{\dagger}_{i}a^{\dagger}_{i}a_{i}a_{i},$ (6)
@@
where the non-Hermitian effective Hamiltonian $\mathcal{H}$ is derived from (3) by adding loss to each site.
$g^{(2)}_{ij}(\tau)=\lim\limits_{t\rightarrow+\infty}\frac{\langle a^{\dagger}_{j}(t)a^{\dagger}_{i}(t+\tau)a_{i}(t+\tau)a_{j}(t)\rangle}{\langle a^{\dagger}_{j}(t)a_{j}(t)\rangle\langle a^{\dagger}_{i}(t+\tau)a_{i}(t+\tau)\rangle},$ (8)
where the expectation values are taken over the vacuum state. For photon blockade, we are only concerned with self-correlation ($i=j$), but we leave $i,j$ arbitrary for now. To derive (8), we develop the perturbative steady-state solution for (5) to first order in $\alpha$. The derivation is presented in the Supplemental Materials [55], and uses the eigendecomposition of the coupling matrix $[J_{ij}]|\varphi_{n}\rangle=\epsilon_{n}|\varphi_{n}\rangle$, where each $|\varphi_{n}\rangle$ is a single-particle eigenmode and $\epsilon_{n}$ is the corresponding eigenenergy relative to the driving frequency. We also define the two-photon basis $|\varphi_{mn}\rangle=|\varphi_{m}\rangle\otimes|\varphi_{n}\rangle$. The result is
$g^{(2)}_{ij}(\tau)$ $=\left|1-e^{-iz\tau}\,\frac{2\alpha}{G_{id}G_{jd}}\sum_{k}G_{kd}^{2}G^{(2)}_{ijkk}(\tau)\right|^{2},$ (9)
$G_{ij}$ $\equiv\sum_{n}\frac{\langle i|\varphi_{n}\rangle\langle\varphi_{n}|j\rangle}{z+\epsilon_{n}},$ (10)
$G^{(2)}_{ijkk}(\tau)$ $\equiv\sum_{mn}\frac{\langle i,j|\varphi_{mn}\rangle\langle\varphi_{mn}|k,k\rangle}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (11)
Here, $|i\rangle\equiv a_{i}^{\dagger}|\varnothing\rangle$ is the state with one photon in cavity $i$; $|i,j\rangle\equiv|i\rangle\otimes|j\rangle$; and $G\equiv\sum_{n}|\varphi_{n}\rangle\langle\varphi_{n}|/(z+\epsilon_{n})$ is the single-particle Green’s function for the linear system.
The time-dependent correlation function (9) applies to any configuration of coupled cavities with weak nonlinearity and weak driving, and has not been previously reported to our knowledge. From it, we see that the $\tau$-dependence of the correlation has two sources. The first is the $e^{-iz\tau}$ factor, which simply accounts for the intrinsic frequency detuning and decay of the cavities; if $\Delta$ is small (i.e., the cavities are driven close to resonance), this reduces to $\approx e^{-\gamma\tau/2}$. This $\tau$-dependence is consistent with the conventional photon blockade result $g^{(2)}(\tau)\approx(1-e^{-\gamma\tau/2})^{2}$ [55] (note, however, that conventional blockade uses strong nonlinearity, $\alpha\gg\gamma$).
The remaining $\tau$-dependence in (9) is from $G^{(2)}_{ijkk}(\tau)$, representing interferometric effects. In Eq. (11), this is written as a superposition of Rabi-like oscillations with eigenfrequencies $\epsilon_{n}$. The photon antibunching time-window is therefore determined by $\max\{|\epsilon_{n}|\}$.
@@
UPB with two cavities—To illustrate the use of this formalism, consider the two-cavity UPB model with $J_{12}=J_{21}=J\in\mathbb{R}^{+}$, for which the eigenmodes are $|\varphi_{1,2}\rangle=(|1\rangle\pm|2\rangle)/\sqrt{2}$, with $\epsilon_{1,2}=\pm J$. Plugging these into Eqs. (9)–(11), we find that the antibunching condition $g_{11}^{(2)}(0)=0$ is satisfied if and only if $z^{3}\approx-\alpha J^{2}/2$ [55]. In Fig. 1(a), the three roots for $z$ are plotted as blue circles. They are symmetrically distributed around the origin (marked by a red cross), the significance of which will be explained shortly. We focus on the root with positive cavity loss, for which
@@
Unlike conventional photon blockade, this is achievable for $\alpha\ll\gamma$, but only in the strong-coupling regime (1) [30, 32]. Moreover, from Eqs. (9)–(11) we can derive
$g_{11}^{(2)}(\tau)$ $\approx\left(1-e^{-\frac{\gamma}{2}\tau}\cos(\Delta\tau)\cos(J\tau)\right)^{2},$ (13)
yielding a short time window of $\delta\tau\sim 1/J$.
Single-particle dark state—The condition for photon antibunching in a designated signal cavity $s$ is $g^{(2)}_{ss}(0)=0$. Using Eq. (9), this can be re-expressed as
$f_{ss}(z)\equiv\frac{G_{sd}^{2}-2\alpha\sum_{k}G_{kd}^{2}G^{(2)}_{sskk}(0)}{G_{sd}^{2}}=0.$ (14)
The Green’s functions, and hence $f_{ss}(z)$, are meromorphic functions in the complex $z$ plane. We claim that for small $\alpha$, the zeros of $f_{ss}(z)$ lie close to the zeros of $G_{sd}(z)$. To see why, observe that $f_{ss}(z)=1$ over the whole $z$-plane when $\alpha=0$. For $\alpha>0$, $f_{ss}$ develops poles at discrete points $\{z_{n}\}$ which are the zeros of $G_{sd}(z)$. By Cauchy’s argument principle, the net number of zeros and poles (counted with their multiplicities and orders) near each $z_{n}$ is conserved, so one or more zeros of $f_{ss}(z)$ must emerge continuously from $z_{n}$. This completes the argument.
The zeros of $G_{sd}(z)$, which we call SPDSs, are instances where the linear version of the system experiences complete destructive interference in the signal cavity. Motivated by the above discussion, we seek a system with an SPDS in the $\gamma\gg\max\{|\epsilon_{n}|\}$ regime. Since $\max\{|\epsilon_{n}|\}$ is the spectral radius of $[J_{ij}]$, which for an $N\times N$ matrix satisfies $\max\{J_{ij}\}\leq\max\{\absolutevalue{\epsilon_{n}}\}\leq N\max\{J_{ij}\}$, we can reformulate the criterion as
[…]
Long-Lived Photon Blockade—Fig. 1(b) shows an exemplary four-cavity setup that can satisfy (15). The cavities are arranged in a ring with couplings $J_{12},J_{23},J_{34},J_{41}\in\mathbb{R}^{+}$. The driven cavity is $d=1$, and the signal cavity is $s=2$.
To understand intuitively how the SPDS arises, we use a Dyson series in the variable $1/z$, which is the Green’s function for an isolated cavity. Each term in the series can be interpreted as a photon trajectory starting from $d$ and ending at $s$ [56]; since we intend $\gamma$ to be large relative to the coupling rates, long trajectories that hop multiple times are negligible. The Dyson series for $G_{21}$ to order $1/z^{4}$ contains the following five trajectories:
Trajectory Dyson series term
$1\mapsto 2$ $\frac{1}{z}\cdot J_{12}\cdot\frac{1}{z}$
Equating the sum of these terms to zero, we obtain [55]
$z=\pm i\sqrt{J_{41}^{2}+J_{12}^{2}+J_{23}^{2}+\frac{J_{41}J_{23}J_{34}}{J_{12}}}.$ (16)
By choosing $J_{12}\ll J_{14},J_{23},J_{34}$, we can get a root with $\gamma=-2\mathrm{Im}(z)\approx 2(J_{41}J_{23}J_{34}/J_{12})^{1/2}$, satisfying Eq. (15).
This four-cavity setup appears to be a minimal configuration for achieving LLPB. Applying the above Dyson series analysis (up to second-lowest order in $1/z$) to two- and three-cavity systems, we find no configurations supporting an SPDS at large $\gamma$. For two cavities, the Dyson series yields $z=\pm iJ_{12}$, whereas for three cavities it gives $z=-J_{12}-J_{13}$ or $z=-J_{13}J_{23}/J_{12}$; none of these solutions satisfy the criterion (15) [55].
We now adopt a mirror-symmetric configuration $J_{12}=J^{\prime}/k$, $J_{14}=J_{23}=J$, and $J_{34}=J^{\prime}$. The above condition for the SPDS can then be met by choosing $J\gtrsim J^{\prime}$ and $k\gg 1$. One potential advantage is that this can be realized with the structure shown in Fig. 1(c), consisting of two ring resonators each hosting a pair of weakly-coupled clockwise and counterclockwise modes [57, 58, 59, 60]. In the Supplemental Materials, we present further details about this coupled-ring setup, including full-wave simulations showing how to achieve the desired model parameters [55]. If we further take $J\gg J^{\prime}$, we can obtain the closed-form approximation
$f_{22}(z)$ $\approx\frac{(z-z_{0}-\delta z)(z-z_{0}+\delta z)}{(z-z_{0})^{2}},$ (17)
$z_{0}\approx$ $-i\sqrt{k}J,\hskip 10.0pt\delta z\approx\pm{\frac{\sqrt{38}}{16}}\sqrt{\alpha\gamma}e^{-i\pi/4}.$ (18)
The SPDS occurs at $z=z_{0}$, with $|\mathrm{Im}(z_{0})|=\sqrt{k}J\gg J$, thus satisfying Eq. (15). This point, along with the nearby zeros of $f_{22}(z)$—where LLPB is predicted to occur—are depicted in Fig. 1(d).
@@
In the upper panel of Fig. 1(e), we plot $g_{22}^{(2)}(0)$ against $\Delta$ and $\gamma$, calculated from the crude approximation Eqs. (17)–(18) using the system parameters stated in the figure caption. Note that the vertical axis is flipped relative to Fig. 1(d). We observe strong local minima, consistent with the above predictions. For comparison, the lower panel of Fig. 1(e) shows results from wavefunction Monte Carlo (WFMC) [61, 62, 63, 64] simulations that directly solve the Lindblad equation (4) for the same cavity parameters, using $F_{d}=10^{-5}$ and Fock cutoffs chosen per site to balance accuracy and efficiency. Details about the simulations may be found in the Supplemental Materials [55]. One notable difference is that the WFMC simulations have $g_{22}^{(2)}(0)\ll 1$ but not exactly zero; this residual value increases with $F_{d}$, alongside an increase in the mean photon occupation number in the signal cavity [55].